OpenStudy (anonymous):

Can someone please check my work? This should come as an easy problem to most folks! FAN AND MEDAL

OpenStudy (anonymous):

The graph of f(x) = x2 is shifted 3 units to the right to obtain the graph of g(x). Which of the following equations best describes g(x)? g(x) = (x + 3)2 g(x) = x2 - 3 g(x) = x2 + 3 g(x) = (x - 3)2

OpenStudy (anonymous):

I'll help

OpenStudy (anonymous):

The twos are exponents btw

OpenStudy (anonymous):

Never mind

OpenStudy (mony01):

its the third one

OpenStudy (anonymous):

That's what I out. Just wanted to double check. Thanks!!

OpenStudy (anonymous):

put**